OpenStudy (anonymous):

Complete the square x^2 + 8x + 11.

OpenStudy (anonymous):

Half the middle term, then square it

OpenStudy (anonymous):

=x^2+8x+11+5-5 =(x^2+8x+16)-5 =(x+4)^2-5

OpenStudy (anonymous):

well thats tricky

OpenStudy (anonymous):

Would you mind if I actually put my work here and you look at what I've done? I should've done that in the first place. Because I don't quite understand what you put there.

OpenStudy (anonymous):

That's actually great idea

OpenStudy (anonymous):

x^2 + 8x + 11 16 + x^2 + 8x + 11 = 0 - 11 x^2 + 8x + 16 = 5 (x + 4)^2 = 5 sqrt(x + 4)^2) = sqrt(5) x + 4 = sqrt(5)

OpenStudy (anonymous):

Your working is correct, firstyoufall. First, it is usually considered an equation.\[x ^{2}+8x +11=0\]Also, you should stop at\[(x +4)^{2}=5\]unless there is some reason to go further.

OpenStudy (anonymous):

You don't have to seperate the (x + 4)^2, like to get rid of the exponent? I'm using the complete the square method for finding x-intercepts.

OpenStudy (anonymous):

OK, if that is what you are doing, go ahead. Usually this technique is used to actually get the square, for example to put an equation in the standard form of an ellipse.

OpenStudy (anonymous):

If you're sure that's the right answer. I guess I just don't know how to put the 2 intercepts in square root form. But thank you for your help.

OpenStudy (anonymous):

What are you asking? That wasn't part of your question. Do you want to solve for x?

OpenStudy (anonymous):

I know, that's why I didn't ask, ha. You don't have to answer anything else. (I was supposed to find the sqrt(5), but they were supposed to stay in radical form. It didn't make sense to me.)